# Rendering Wireless Environments Useful for Gradient Estimators: A Zero-Order Stochastic Federated Learning Method

### v1

@@ C-B Proof of Theorem 1: Convergence Analysis
$\begin{split}&\sum_{k=0}^{\infty}\alpha_{k+1}\gamma_{k+1}\|\nabla F(\theta_{k+1})\|^{2}\\ \geq&(\rho-\varepsilon)^{2}\sum_{k=0}^{\infty}\alpha_{k+1}\gamma_{k+1}-2(\rho-\varepsilon)L\sqrt{c}\sum_{k=0}^{\infty}\alpha_{k+1}\gamma_{k+1}\alpha_{k}\\ &+L^{2}c\sum_{k=0}^{\infty}\alpha_{k+1}\gamma_{k+1}\alpha_{k}^{2}\\ =&+\infty,\end{split}$ (20)
as the first series diverges, and the second and the third converge by Assumption 5. This implies that the series $\sum_{k}\alpha_{k}\gamma_{k}\|\nabla F(\theta_{k})\|^{2}$ diverges. This is a contradiction as this series converges almost surely by (18). Therefore, hypothesis (H) cannot be true and $\|\nabla F(\theta_{k})\|$ converges to zero almost surely.
@@ C-C Proof of Theorem 2: Convergence Rate
Starting again from the $L$-smoothness in Lemma 3 and the algorithm update step $\theta_{k+1}=\theta_{k}-\alpha_{k}g_{k}$, we have
$\begin{split}F(\theta_{k+1})&\leq F(\theta_{k})-\alpha_{k}\langle\nabla F(\theta_{k}),g_{k}\rangle+\frac{\alpha_{k}^{2}L}{2}\|g_{k}\|^{2}.\end{split}$ (21)
Taking the conditional expectation given $\mathcal{H}_{k}$,
$\begin{split}&F(\theta_{k+1})\\ \leq&F(\theta_{k})-c_{1}\alpha_{k}\gamma_{k}\langle\nabla F(\theta_{k}),\nabla F(\theta_{k})+b_{k}\rangle+\frac{\alpha_{k}^{2}Lc_{2}}{2}\\ =&F(\theta_{k})-c_{1}\alpha_{k}\gamma_{k}\|\nabla F(\theta_{k})\|^{2}-c_{1}\alpha_{k}\gamma_{k}\langle\nabla F(\theta_{k}),b_{k}\rangle\\ &+\frac{\alpha_{k}^{2}Lc_{2}}{2}\\ \end{split}$
$\begin{split}\overset{(a)}{\leq}&F(\theta_{k})-c_{1}\alpha_{k}\gamma_{k}\|\nabla F(\theta_{k})\|^{2}+\frac{c_{1}\alpha_{k}\gamma_{k}}{2}\|\nabla F(\theta_{k})\|^{2}\\ &+\frac{c_{1}\alpha_{k}\gamma_{k}}{2}\|b_{k}\|^{2}+\frac{\alpha_{k}^{2}Lc_{2}}{2}\\ =&F(\theta_{k})-\frac{c_{1}\alpha_{k}\gamma_{k}}{2}\|\nabla F(\theta_{k})\|^{2}+\frac{c_{1}\alpha_{k}\gamma_{k}}{2}\|b_{k}\|^{2}+\frac{\alpha_{k}^{2}Lc_{2}}{2}\end{split}$ (22)
where $(a)$ is by $-\langle a,b\rangle\leq\frac{1}{2}\|a\|^{2}+\frac{1}{2}\|b\|^{2}$
By considering a large value $K>0$ and taking the telescoping sum of (22), we get
$\begin{split}&\mathbb{E}[F(\theta_{K+1})|\mathcal{H}_{K}]\\ \leq&F(\theta_{0})-\frac{c_{1}}{2}\sum_{k=0}^{K}\alpha_{k}\gamma_{k}\|\nabla F(\theta_{k})\|^{2}+\frac{c_{1}}{2}\sum_{k=0}^{K}\alpha_{k}\gamma_{k}\|b_{k}\|^{2}\\ &+\frac{Lc_{2}}{2}\sum_{k=0}^{K}\alpha_{k}^{2}.\end{split}$ (23)
Given the assumption that $F(\theta^{*})=\min_{\theta\in\mathbb{R}^{d}}F(\theta)$ exists, we know that $\delta_{k}=F(\theta_{k})-F(\theta^{*})\geq 0$. Then,
$\begin{split}0\leq\mathbb{E}[\delta_{K+1}|\mathcal{H}_{K}]\leq&\delta_{0}-\frac{c_{1}}{2}\sum_{k=0}^{K}\alpha_{k}\gamma_{k}\|\nabla F(\theta_{k})\|^{2}\\ &+\frac{c_{1}c_{3}}{2}\sum_{k=0}^{K}\alpha_{k}\gamma_{k}^{3}+\frac{Lc_{2}}{2}\sum_{k=0}^{K}\alpha_{k}^{2}.\end{split}$ (24)
Finally,
$\begin{split}\sum_{k}\alpha_{k}\gamma_{k}\mathbb{E}[\|\nabla F(\theta_{k})\|^{2}]&\leq\frac{2}{c_{1}c_{3}}\delta_{0}+\sum_{k}\alpha_{k}\gamma_{k}^{3}+\frac{Lc_{2}}{c_{1}}\sum_{k}\alpha_{k}^{2}\\ \end{split}$ (25)
Let $\alpha_{k}$ and $\gamma_{k}$ have the forms given in Example 2. We know that, $\forall K>0$,
@@
Similarly, $\sum_{k=0}^{K}\alpha_{k}^{2}\leq\alpha_{0}^{2}\big{(}\frac{2\upsilon_{1}}{2\upsilon_{1}-1}\big{)}$. Next, when further $10<\upsilon_{1}+\upsilon_{2}<$ 1,
@@
Thus, making use of inequality (25)
$\frac{\sum_{k}\alpha_{k}\gamma_{k}\mathbb{E}[\|\nabla F(\theta_{k})\|^{2}]}{\sum_{k}\alpha_{k}\gamma_{k}}\leq\frac{A(1-\upsilon_{1}-\upsilon_{2})}{(K+2)^{1-\upsilon_{1}-\upsilon_{2}}-1}$ (28)
with $A=\frac{2}{c_{1}\alpha_{0}\gamma_{0}}\delta_{0}+c_{3}^{2}\gamma_{0}^{2}\big{(}\frac{\upsilon_{1}+3\upsilon_{2}}{\upsilon_{1}+3\upsilon_{2}-1}\big{)}+\frac{Lc_{2}\alpha_{0}}{c_{1}\gamma_{0}}\big{(}\frac{2\upsilon_{1}}{2\upsilon_{1}-1}\big{)}$
Appendix D Non-Symmetrical Channels
Assuming non-symmetrical channels with $\mathbb{E}[h_{i,k}]=\mu_{h}$ and $\sigma_{h}^{2}=\mathbb{E}[h_{i,k}^{2}]-\mu_{h}^{2}$, $\forall i,\forall k$, the one-point gradient estimate becomes
$g_{k}=\Phi_{k}\sum_{i=1}^{N}\bigg{[}h_{i,k+1}\tilde{f}_{i}\big{(}\theta_{k}+\gamma_{k}\Phi_{k},S_{i,k+1}\big{)}+n_{i,k+1}\bigg{]}.$ (29)
And 1P-ZOFL is updated to Algorithm 2.
Algorithm 2 The 1P-ZOFL algorithm with non-symmetrical channels
Input: Initial model $\theta_{0}\in\mathbb{R}^{d}$, the initial step-sizes $\alpha_{0}$ and $\gamma_{0}$, and the channels’ standard deviation $\sigma_{h}$
1:  for $k=0,2,4,...$ do
2:     The server broadcasts $\theta_{k}+\gamma_{k}\Phi_{k}$
3:     The server receives $\sum_{i=1}^{N}h_{i,k+1}\tilde{f}_{i}\big{(}\theta_{k}+\gamma_{k}\Phi_{k},S_{i,k+1}\big{)}+n_{i,k+1}$
4:     The server multiplies the received scalar sum by $\Phi_{k}$ to assemble $g_{k}$ in (29)
5:     The server updates $\theta_{k+1}=\theta_{k}-\alpha_{k}g_{k}$
6:  end for
We then analyze the properties of our modified gradient estimate:
$\begin{split}&\mathbb{E}[g_{k}|\mathcal{H}_{k}]\\ =&\mathbb{E}\Big{[}\Phi_{k}\sum_{i=1}^{N}\Big{(}h_{i,k+1}\tilde{f}_{i}\big{(}\theta_{k}+\gamma_{k}\Phi_{k},S_{i,k+1}\big{)}+n_{i,k+1}\Big{)}\Big{|}\mathcal{H}_{k}\Big{]}\\ =&\mathbb{E}\Big{[}\Phi_{k}\sum_{i=1}^{N}\mu_{h}\tilde{F}_{i}\big{(}\theta_{k}+\gamma_{k}\Phi_{k}\big{)}\Big{|}\mathcal{H}_{k}\Big{]}\\ =&\mu_{h}\gamma_{k}\sum_{i=1}^{N}\mathbb{E}\Big{[}\Phi_{k}\Phi_{k}^{T}\nabla\tilde{F}_{i}(\theta_{k})+\gamma_{k}\Phi_{k}\Phi_{k}^{T}\nabla^{2}\tilde{F}_{i}(\breve{\theta}_{k})\Phi_{k}\Big{|}\mathcal{H}_{k}\Big{]}\\ =&c_{1}\gamma_{k}(\nabla F(\theta_{k})+b_{k}),\\ \end{split}$ (30)
with $c_{1}=\frac{\mu_{h}\alpha_{2}}{\sigma_{h}^{2}}$ and $b_{k}=\frac{\gamma_{k}}{\alpha_{2}}\mathbb{E}\Big{[}\Phi_{k}\Phi_{k}^{T}\nabla^{2}F_{i}(\breve{\theta}_{k})\Phi_{k}\Big{|}\mathcal{H}_{k}\Big{]}$.
Then, $\|b_{k}\|\leq\frac{\gamma_{k}}{\alpha_{2}}\alpha_{3}^{3}\alpha_{1}$, now $c_{3}=\frac{\alpha_{3}^{3}\alpha_{1}}{\alpha_{2}}$ and $\|b_{k}\|\leq c_{3}\gamma_{k}$.
Let $\theta_{k}^{\prime}=\theta_{k}+\gamma_{k}\Phi_{k}$,
$\begin{split}&\mathbb{E}[\|g_{k}\|^{2}|\mathcal{H}_{k}]\\ =&\mathbb{E}\bigg{[}\|\Phi_{k}\|^{2}\Big{(}\sum_{i=1}^{N}h_{i,k+1}\tilde{f}_{i}\bigg{(}\theta_{k}^{\prime},S_{i,k+1}\bigg{)}+n_{i,k+1}\Big{)}^{2}\Big{|}\mathcal{H}_{k}\bigg{]}\\ =&\alpha_{3}^{2}N\sum_{i=1}^{N}\mathbb{E}\bigg{[}h_{i,k+1}^{2}\tilde{f}_{i}^{2}\bigg{(}\theta_{k}^{\prime},S_{i,k+1}\bigg{)}+n_{i,k+1}^{2}\Big{|}\mathcal{H}_{k}\bigg{]}\\ \leq&\alpha_{3}^{2}N\sum_{i=1}^{N}\mathbb{E}\bigg{[}\frac{h_{i,k+1}^{2}}{\sigma_{h}^{4}}\Big{(}\|f_{i}(0,S_{i,k+1})\|+L_{S_{i,k+1}}\|\theta_{k}^{\prime}\|\Big{)}^{2}\Big{|}\mathcal{H}_{k}\bigg{]}\\ &+N^{2}\alpha_{3}^{2}\sigma_{n}^{2}\\ \end{split}$
$\begin{split}\leq&3\alpha_{3}^{2}N\sum_{i=1}^{N}\mathbb{E}\bigg{[}\frac{h_{i,k+1}^{2}}{\sigma_{h}^{4}}\Big{(}\|f_{i}(0,S_{i,k+1})\|^{2}+L_{S_{i,k+1}}^{2}\|\theta_{k}\|^{2}\\ &\hskip 85.35826pt+\alpha_{3}^{2}L_{S_{i,k+1}}^{2}\gamma_{k}^{2}\Big{)}\Big{|}\mathcal{H}_{k}\bigg{]}+N^{2}\alpha_{3}^{2}\sigma_{n}^{2}\\ =&\frac{3N^{2}\alpha_{3}^{2}(\sigma_{h}^{2}+\mu_{h}^{2})}{\sigma_{h}^{2}}\Big{(}\mu_{S}+L_{S}\|\theta_{k}\|^{2}+\alpha_{3}^{2}L_{S}\gamma_{k}^{2}\Big{)}+N^{2}\alpha_{3}^{2}\sigma_{n}^{2}\\ :=&c_{2}.\end{split}$ (31)
Appendix E Experimental results details
E-A Autoencoder
We compress images from the MNIST dataset using a lossy autoencoder with encoder-decoder architecture. The encoder reduces the input size from 784 to 10 through three linear layers with dimensions 784-512-128-10. The first two layers have ELU activation functions. The decoder reverses this process, going from 10 to 784 through three linear layers (10-128-512-784), with ELU activation in the first two layers and a sigmoid activation in the last layer to ensure output pixel intensities between 0 and 1. Training is done on the MNIST dataset for 10 epochs, using mean squared error loss and the Adam optimizer.
E-B Performance of 1P-ZOFL vs SNR

### v2

@@ C-B Proof of Theorem 1: Convergence Analysis
$\begin{split}\sum_{k=0}^{\infty}\alpha_{k+1}\gamma_{k+1}\|\nabla F(\theta_{k+1})\|^{2}&\geq(\rho-\varepsilon)^{2}\sum_{k=0}^{\infty}\alpha_{k+1}\gamma_{k+1}\\ &-2(\rho-\varepsilon)L\sqrt{c}\sum_{k=0}^{\infty}\alpha_{k+1}\gamma_{k+1}\alpha_{k}\\ &+L^{2}c\sum_{k=0}^{\infty}\alpha_{k+1}\gamma_{k+1}\alpha_{k}^{2}\\ &>+\infty.\end{split}$
On the R.H.S., the first series diverges, and the second and the third converge by Assumption 5. This implies that $\sum_{k}\alpha_{k}\gamma_{k}\|\nabla F(\theta_{k})\|^{2}$ diverges. This is a contradiction as this series converges almost surely by (15). Therefore, hypothesis (H) cannot be true and $\|\nabla F(\theta_{k})\|$ converges to zero almost surely.
@@ C-C Proof of Theorem 2: Convergence Rate
Taking the expectation on both sides of (14) results in $\mathbb{E}[\langle\nabla F(\theta_{k}),e_{k}\rangle]=\mathbb{E}_{\mathcal{H}_{k}}\Big{[}\mathbb{E}\big{[}\langle\nabla F(\theta_{k}),e_{k}\rangle\big{|}\mathcal{H}_{k}\big{]}\Big{]}=0$. Then, substituting by the upper bound of Lemma 2,
$\begin{split}\sum_{k=0}^{K}\alpha_{k}\gamma_{k}\mathbb{E}[\|\nabla F(\theta_{k})\|^{2}]&\leq\frac{2}{c_{1}}\delta_{0}+c_{3}\sum_{k=0}^{K}\alpha_{k}\gamma_{k}^{3}+\frac{Lc_{2}}{c_{1}}\sum_{k=0}^{K}\alpha_{k}^{2}\\ \end{split}$ (16)
@@
Thus, making use of inequality (16),
$\frac{\sum_{k}\alpha_{k}\gamma_{k}\mathbb{E}[\|\nabla F(\theta_{k})\|^{2}]}{\sum_{k}\alpha_{k}\gamma_{k}}\leq\frac{A(1-\upsilon_{1}-\upsilon_{2})}{(K+2)^{1-\upsilon_{1}-\upsilon_{2}}-1},$ (17)
with $A=\frac{2}{c_{1}\alpha_{0}\gamma_{0}}\delta_{0}+c_{3}^{2}\gamma_{0}^{2}\big{(}\frac{\upsilon_{1}+3\upsilon_{2}}{\upsilon_{1}+3\upsilon_{2}-1}\big{)}+\frac{Lc_{2}\alpha_{0}}{c_{1}\gamma_{0}}\big{(}\frac{2\upsilon_{1}}{2\upsilon_{1}-1}\big{)}$.
We see that $\upsilon_{1}=\frac{1}{2}$ and $\upsilon_{2}=\frac{1}{6}$ are the optimal choice for the time-varying part of the rate, i.e., for $O(\frac{1}{\sqrt[3]{K}})$. However, to prevent the constant part from growing too large, we choose slightly larger exponents of $\upsilon_{1}=\frac{1}{2}+\frac{\epsilon}{2}$ and $\upsilon_{2}=\frac{1}{6}+\frac{\epsilon}{2}$, where $\epsilon$ is a small strictly positive value. This will result in a rate of $O\big{(}\frac{1}{K^{\frac{1}{3}-\epsilon}}\big{)}$.
Appendix D Performance of 1P-ZOFL vs SNR